# Verifiable blind observable estimation: A composably secure protocol for near-term quantum advantage tasks

### v1

@@ II.4 Concrete construction of the SDOE resource
$\begin{split}&\max_{m}\Pr[Z\geq(kw+\gamma_{2})N_{c},Y\leq w]\\ &=\max\left\{\max_{m\leq m_{0}}\Pr[Z\geq(kw+\gamma_{2})N_{c},Y\leq w],\right.\\ &\quad\quad\quad\quad\left.\max_{m>m_{0}}\Pr[Z(\geq kw+\gamma_{2})N_{c},Y\leq w]\right\}\\ &<\max_{m\leq m_{0}}\Pr[Z\geq(kw+\gamma_{2})N_{c}]+\max_{m>m_{0}}\Pr[Y\leq w].\end{split}$ (5)
@@
Within this setting, for a given value of $m$, the tail bound for the hypergeometric distribution gives
@@
$\begin{split}&\Pr[Z\geq(kw+\gamma_{2})N_{c}]\\ &\leq\exp\left(-2N_{c}\left((kw+\gamma_{2})-\frac{m}{N_{c}+N_{t}}\right)^{2}\right),\end{split}$ (6)
so that we obtain
$\begin{split}&\max_{m}\Pr[Z\geq(kw+\gamma_{2})N_{c},Y\leq w]\\ &=\max\left\{\max_{m\leq m_{0}}\Pr[Z\geq(kw+\gamma_{2})N_{c},Y\leq w],\right.\\ &\quad\quad\quad\quad\left.\max_{m>m_{0}}\Pr[Z(\geq kw+\gamma_{2})N_{c},Y\leq w]\right\}\\ &<\max_{m\leq m_{0}}\Pr[Z\geq(kw+\gamma_{2})N_{c}]+\max_{m>m_{0}}\Pr[Y\leq w].\end{split}$ (7)
@@
To bound $\Pr[Y\leq w]$ with $m\geq kw+\gamma_{2}/2$, recall that test rounds are defined as in RVBQC (Protocol 5) and that a given deviation on a test round is detected with probability at least $\frac{1}{k}$, where $k$ is the chromatic number of $G$. This means that conditioned on $X$, the number of failed test rounds $Y$ is lower bounded in the usual stochastic order by a binomial distribution with $X$ samples and average $1/k$. This implies that, with $m>m_{0}$
@@
$\begin{split}&\Pr[Y\leq w]\\ &=\Pr\left[Y\leq w,X\leq N_{t}\left(\frac{m_{0}}{N_{c}+N_{t}}-\frac{\gamma_{2}}{4}\right)\right]\\ &\quad+\Pr\left[Y\leq w,X>N_{t}\left(\frac{m_{0}}{N_{c}+N_{t}}-\frac{\gamma_{2}}{4}\right)\right]\\ &<\Pr\left[Y\leq w,X\leq N_{t}\left(kw+\frac{\gamma_{2}}{4}\right)\right]\\ &\quad+\Pr\left[Y\leq w,X>N_{t}\left(kw+\frac{\gamma_{2}}{4}\right)\right]\\ &<\Pr\left[X\leq N_{t}\left(kw+\frac{\gamma_{2}}{4}\right)\right]\\ &\quad+\Pr\left[Y\leq w|X=N_{t}\left(kw+\frac{\gamma_{2}}{4}\right)\right]\\ &<\exp\left(-N_{t}\frac{\gamma_{2}^{2}}{2}\right)+\exp\left(-\frac{\gamma_{2}^{2}}{8k^{2}}\frac{N_{c}}{kw+\gamma_{2}/4}\right).\end{split}$ (8)
@@
Because both Equations 6 and 8 provide bounds that are negligible in $N_{c}$ and $N_{t}$, this shows that the distinguishing advantage provided by the ideal SDOE resource rejecting more often than the VBOE protocol is indeed negligible in $N_{c}$ for a fixed ratio $N_{c}/N_{t}$. Hence, we conclude that VBOE is constructing SDOE with negligible error in $N_{c}$ provided that $kw$ is below and bounded away from $\epsilon$ by a constant, so the positive constants $\gamma_{1}$ and $\gamma_{2}$ can be set such that $\gamma_{1}+(kw+\gamma_{2})<\epsilon$. ∎
@@ III Discussion
We have introduced the verifiable blind observable estimation (VBOE) protocol, extending the framework of verifiable blind delegated quantum computation protocols [13, 14] to effectively and securely delegate to the server the observable estimation tasks that underpin various quantum tasks. We obtain Theorem 1 that guarantees the composable security with negligible error of the VBOE protocol.
@@
Our primary contribution is a fundamental enhancement of the available toolkit for verification of quantum computations in the AC framework. It enables the rigorous security analysis of algorithms relying on observable estimation. We achieve this by introducing a new conceptual tool: an ideal resource (SDOE) that formally incorporates bounded estimation error, reflecting the statistical nature of near-term algorithms. This extension is not merely a technicality; it provides the only known path to achieving efficient, i.e. polynomial-time, verification with an exponentially small soundness error for this broad and practical class of computations. Our protocol, VBOE, is the first to realise this new paradigm.
The second contribution lies in the limited overhead of our protocol. The conventional approach to securely search for the accurate estimate of the expectation value with RVBQC would be to decompose the estimation task into a series of decision problems with a bisection method [31]. However, converting an estimation task into a decision problem requires additional circuit overhead to implement a POVM that extracts the information whether the target value is in the partitioned region. This overhead would be particularly significant for an MBQC model with a fixed MBQC pattern. Besides, this bisection approach repetitively calls RVBQC as a subroutine, which accordingly increases the distinguishing advantage in total. Therefore, the RVBQC protocol is not optimal for estimation problems.
In contrast, the VBOE protocol keeps the same structure as the original circuit. By directly averaging over the outputs of computation rounds, VBOE offloads the circuit overhead to the sample overhead. This makes the verification of observable estimation much more practical, particularly under the limited availability of quantum resources.
The applicability of the VBOE protocol opens several promising directions for near-term applications. Given that VBOE returns expectation values, a key open question is the secure integration of quantum error mitigation [32, 33, 34, 35, 36, 37] with quantum verification, which could enhance verification noise-robustness while enabling credible error mitigation. Another compelling direction lies in the synergy between multi-party variants of VBOE and device-efficient near-term technologies such as hybrid tensor networks, which may lead to a secure quantum–classical framework that decomposes large simulation tasks into smaller quantum and classical components [38, 39, 40].
With the rapid progress of quantum hardware, a variety of platforms for delegated quantum computation are becoming available. Notably, the first experimental demonstration of UBQC was achieved on photonic devices [41], highlighting the natural compatibility of MBQC with photonic architectures. More recently, RVBQC has been successfully implemented on trapped-ion devices [22], further demonstrating the feasibility of running verification protocols on real hardware.
@@
Building on these developments, the proposed VBOE protocol can also be executed on available quantum platforms, where practical applications such as observable estimation are within reach. We envisage that applying our method on such platforms would enable end-to-end verification for more meaningful computational tasks, beyond proof-of-principle experiments. This would represent a significant step towards bridging theoretical protocols with practical, device-level implementations of verifiable delegated quantum computation.

### v2

@@ II.4 Concrete construction of the SDOE resource
$\begin{split}&\max_{m}\operatorname{Pr}\left[Z\geq(K\omega+\gamma_{2})N_{c},Y\leq\omega N_{t}\right]\\ &=\max\left\{\max_{m\leq m_{0}}\operatorname{Pr}\left[Z\geq(K\omega+\gamma_{2})N_{c},Y\leq\omega N_{t}\right],\right.\\ &\quad\quad\quad\quad\left.\max_{m>m_{0}}\operatorname{Pr}\left[Z\geq(K\omega+\gamma_{2})N_{c},Y\leq\omega N_{t}\right]\right\}\\ &<\max_{m\leq m_{0}}\operatorname{Pr}\left[Z\geq(K\omega+\gamma_{2})N_{c}\right]+\max_{m>m_{0}}\operatorname{Pr}\left[Y\leq\omega N_{t}\right],\end{split}$ (5)
where $m_{0}=\left(K\omega+\frac{\gamma_{2}}{2}\right)(N_{c}+N_{t})$.
@@
Within this setting, for a given value of $m\leq m_{0}$, the tail bound for the hypergeometric distribution gives
@@
$\begin{split}&\operatorname{Pr}[Z\geq(K\omega+\gamma_{2})N_{c}]\\ &\leq\exp\left(-2\left((K\omega+\gamma_{2})-\frac{m}{N_{c}+N_{t}}\right)^{2}N_{c}\right)\\ &\leq\exp\left(-\frac{\gamma_{2}^{2}}{2}N_{c}\right).\end{split}$ (6)
@@
To bound $\operatorname{Pr}\left[Y\leq\omega N_{t}\right]$ with $m>m_{0}$, recall that test rounds are defined as Protocol 3 in RVBQC and that a given deviation on a test round is detected with probability at least $1/K$, where $K$ is the chromatic number of the underlying graph $G$. This means that conditioned on $X$, the number of failed test rounds $Y$ is lower bounded in the usual stochastic order by a binomial distribution with $X$ samples and average $1/K$. This implies that, with $m>m_{0}$,
@@
$\begin{split}&\operatorname{Pr}\left[Y\leq\omega N_{t}\right]\\ &=\operatorname{Pr}\left[Y\leq\omega N_{t},X\leq N_{t}\left(\frac{m_{0}}{N_{c}+N_{t}}-\frac{\gamma_{2}}{4}\right)\right]\\ &\quad+\operatorname{Pr}\left[Y\leq\omega N_{t},X>N_{t}\left(\frac{m_{0}}{N_{c}+N_{t}}-\frac{\gamma_{2}}{4}\right)\right]\\ &=\operatorname{Pr}\left[Y\leq\omega N_{t},X\leq N_{t}\left(K\omega+\frac{\gamma_{2}}{4}\right)\right]\\ &\quad+\operatorname{Pr}\left[Y\leq\omega N_{t},X>N_{t}\left(K\omega+\frac{\gamma_{2}}{4}\right)\right]\\ &<\operatorname{Pr}\left[X\leq N_{t}\left(K\omega+\frac{\gamma_{2}}{4}\right)\right]\\ &\quad+\operatorname{Pr}\left[Y\leq\omega N_{t}~\Big|~X=N_{t}\left(K\omega+\frac{\gamma_{2}}{4}\right)\right]\\ &<\exp\left(-\frac{\gamma_{2}^{2}}{8}N_{t}\right)+\exp\left(-\frac{\gamma_{2}^{2}}{32K^{2}}N_{t}\right).\end{split}$ (7)
@@
Because both Equations (6) and (7) provide bounds that are negligible in $N_{c}$ and $N_{t}$, this shows that the distinguishing advantage provided by the SDOE resource rejecting more often than the VBOE protocol is indeed negligible in $N_{c}$ for a fixed ratio $N_{c}/N_{t}$. Hence, we conclude that VBOE is constructing SDOE with negligible error in $N_{c}$ provided that $K\omega$ is below and bounded away from $\epsilon$ by a constant, so the positive constants $\gamma_{1}$ and $\gamma_{2}$ can be set such that $\gamma_{1}+2(K\omega+\gamma_{2})<\epsilon$. ∎
@@ III Discussion
We have introduced the Verifiable Blind Observable Estimation (VBOE) protocol, which extends the capabilities of verifiable blind delegated quantum computation to the critical task of observable estimation. Our main result, Theorem 1, establishes that VBOE achieves composable security with exponentially negligible error and polynomial-time execution, addressing a longstanding gap in the verification of near-term quantum algorithms.
III.1 Enabling security in the quantum utility regime
@@
Our primary contribution is the extension of the available toolkit for the verification of quantum computations in the Abstract Cryptography framework. This enables embedding observable estimation in a broader, possibly hybrid, secure computational task while not having to reprove its security in this wider context. This is achieved by introducing a new conceptual tool: an ideal resource (SDOE) that formally incorporates bounded estimation error, reflecting the statistical and possibly noisy nature of near-term algorithms. This extension is not merely a technicality; it provides the only known path to achieving efficient polynomial-time verification with an exponentially small soundness error for this broad and practical class of computations. The proposed VBOE protocol is the first to construct this new resource and to outline a path to realising it in practice.
This unlocks the strongest composable security model for a vast set of near-term applications, whose security analysis was previously out of reach. For instance, in recent works, observable estimation problems, such as out-of-time-order correlations, have been put forward as milestone experiments because their quantum advantage is verifiable using another quantum hardware or natural quantum system, while lying beyond the reach of known classical simulation methods [Abanin2025Observation, MK2025Verifiable]. However, even when verification is attempted using another quantum device, the reliability of the outcome remains conceptually unresolved: without a principled cryptographic framework, the trustworthiness of the devices involved cannot be established a priori. Our framework directly addresses this gap by providing a composable, device-agnostic verification protocol for observable estimation outcomes without requiring trust in any particular quantum prover.
Our framework also provides a formal abstract cryptography security guarantee for recent applications that make use of quantum verification protocols. In particular, Inajetovic et al. [Inajetovic2025Verifiable] introduce a verifiable end-to-end delegated variational quantum algorithm (MB-DVQA) by invoking a variant of RVBQC as a subroutine to ensure the correctness of individual optimisation steps. While their work rigorously establishes the correctness and verifiability of the algorithmic outcome, a composable cryptographic security analysis of the overall construction is left open. Using our result, the MB-DVQA protocol can be analysed at the level of cryptographic resources, relying on the composable security of VBOE for delegated observable estimation tasks that were previously unavailable.
III.2 Low implementation overhead
A key practical advantage of VBOE is its low implementation overhead, which sets it apart from previous approaches. Conventional methods for verifying observable estimation with exponential security, such as those based on RVBQC, would typically need to verify both the acquisition of measurement outcomes and the empirical average. This would either blow up the space overhead if the outcomes are acquired in parallel, or require long-term memory to store the partial sum of outcomes if the outcomes are acquired sequentially. In both cases, we would need a fault-tolerant quantum computer.
In contrast, the VBOE protocol is designed to preserve the structure of the original target computation generating a single measurement outcome per round. By letting the client average over the outputs of the computation rounds, VBOE suppresses the circuit overhead of previous approaches while maintaining the security guarantees. As a result, VBOE offers a significantly more practical verification strategy for observable estimation tasks, particularly in regimes where quantum resources are limited.
@@ III.2 Low implementation overhead
We envisage that applying our method on such platforms would enable end-to-end verification for more meaningful computational tasks, beyond proof-of-principle experiments. Following the first experimental demonstration of UBQC on photonic devices [Barz2012Demonstration], recent work has reported the successful implementation of RVBQC on trapped-ion platforms [drmota2024verifiable]. With the rapid progress of quantum hardware, a growing number of high-quality quantum devices with remote access are becoming available [Hughes2025Trapped-ion, Ransford2025Helios]. Building on these developments, the proposed VBOE protocol can likewise be executed on existing quantum platforms, where more practically relevant applications based on observable estimation are within reach. This represents a significant step toward bridging theoretical protocols and practical, device-level implementations of verifiable delegated quantum computation.
III.3 Compatibility with quantum error mitigation
By its very nature as a protocol for verifying expectation values, VBOE is naturally compatible with near-term and early fault-tolerant error-suppression techniques. A key open question is how to securely integrate quantum error mitigation (QEM) [li2017efficient, temme2017error, endo2018practical, yang2022efficient, endo2021hybrid, cai2023quantum], which aims to recover noise-free expectation values through classical post-processing with limited quantum overhead. Integrating QEM with VBOE has the potential to enhance the noise robustness of the protocol by reducing abort probabilities, while simultaneously enabling credible error mitigation with explicit cryptographic guarantees.
Another promising direction lies in combining multi-party variants of VBOE with device-efficient near-term techniques, such as circuit cutting and hybrid tensor-network methods [peng2020simulating, yuan2021quantum, Harada2025densitymatrix, Yang2025Resource-efficient]. Such a synergy could lead to a secure quantum-classical framework in which large-scale simulation tasks are decomposed into smaller quantum and classical subroutines and distributed across multiple remote parties.